OpenStudy (korosh23):

Physics 12 Question! Circular motion! A car travels around the same circular track at different velocities. which graph best shows the relationship between its centripetal acceleration, ac, and its velocity, v?

OpenStudy (korosh23):

ac or a= V^2/r a= acceleration v= velocity r= radius This is the formula, but I do not understand which graph it is?

rvc (rvc):

since it has different velocities and a=v^2/r even acc is increasing so it has an increasing slope..

OpenStudy (korosh23):

I do not understand why ac is increasing, and why it has an increasing slope? Could you please explain more?